OpenStudy (anonymous):

In your own words explain the difference between the following two expressions:

OpenStudy (anonymous):

\[(-1)^\frac{ 2 }{ 4 }\]

OpenStudy (anonymous):

\[ (-1)^\frac{ 1 }{ 2 }\]

jhonyy9 (jhonyy9):

\[\sqrt[4]{(-1)^2}\]

jhonyy9 (jhonyy9):

the first you can rewrite it hence like above - yes ?

jhonyy9 (jhonyy9):

second wan being \[\sqrt{-1}\]

jhonyy9 (jhonyy9):

yes ?

jhonyy9 (jhonyy9):

do you see now what is the difference ?